OpenStudy (anonymous):

Write a rule for the nth term of the arithmetic sequence. 2/1, 3/2, 4/3, 5/4, . . .

ganeshie8 (ganeshie8):

\(\large \frac{2}{1},\frac{3}{2},\frac{4}{3},\frac{5}{4},...\)

ganeshie8 (ganeshie8):

find the nth term of numerator : 2, 3, 4, 5... find the nth term of denominator : 1, 2, 3, 4...

ganeshie8 (ganeshie8):

divide

OpenStudy (anonymous):

what would the answer be ? :0

ganeshie8 (ganeshie8):

2, 3, 4, 5... can u find nth term of this sequence ?

ganeshie8 (ganeshie8):

each next term is getting added by 1, so the nth term is simply : \(n+1\)

ganeshie8 (ganeshie8):

similarly find the nth term for denominator also : 1, 2, 3, 4...

ganeshie8 (ganeshie8):

nth term is just \(n\) here

ganeshie8 (ganeshie8):

divde the numerator and denominator nth terms for overall sequence : \(\large \frac{n+1}{n}\)

OpenStudy (anonymous):

im so lost :( so the answer is n+1 over N ??